answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Why is the Normal Distribution called normal?

Suppose you could call it the Gaussian Distribution or the Laplace-Gauss (not to be confused with the Laplace distribution which takes an absolute difference from the mean rather than a squared error)... however the Brits had no one to name this distribution after (not the German and French names) and because it is the ubiquitous distribution they just called it... well the NORMAL!!

A bag contain 4 green 6 black 7 white balls a ball is drawn at random what is the probability that it is a green or a black ball?

This is a law of addition probability which states that the probability of A or B equals the probability of A plus the probability of B minus the probability of A and B. Written in mathematical terms, the equation is: P(AorB) = P(A) + P(B) - P(AnB) where P(AnB) = 0 (since you can not pull out a green and black ball at the same time). Let P(A) = Probability of drawing the green ball & let P(B) = Probability of drawing the black ball. Total outcomes is 17. So, P(A) = 4/17 & P(B) = 6/17. Therefore P(green or black) = 4/17 + 6/17 = 10/17.

Two coins and one six sided cube are tossed together What is the probability of getting 2 heads and a four?

Probability is defined as the number of ways an outcome can occur divided by the number of possible outcomes. For the coins, there are 4 outcomes (HH, HT, TH, TT). On the cube, there are 6 possible outcomes. The total number of outcomes is then 4*6 = 24. Since there is only 1 way to obtain HH, look at the cube outcomes. With the HH outcome, the cube would need to fall on a 4. So, there is only 1 way a HH4 can occur. Therefore the probability of getting 2 heads and a four is 1/24 or 0.04167.

How many cards must you draw from a regular deck of 52 playing cards to be sure at least 2 are from the same suit?

5. Assuming the first four are all different suits, the 5th card must be a duplicate suit. If any prior to the 5th card is not a new suit, it is garunteed to be a duplicate of a prior suit.

An event that will definitely happen is an event with what percent probability?

An event that will definitely happen is an event with 100% probability.

How do you find the area to the right of the z-score?

Charts typically show and list the area to the left of the Z-Score value. To find the area to the right, just subtract the Z-Score value from 1; e.g. if the Z-Score value is .75 then take 1-.75 = .25.

Three fair coins are tossed determine the probability of the following event exactly two coins land on heads?

The sample space is 23 or 8; which can be listed out as: HHH, HHT, HTH, HTT, THH, THT, TTH, TTT. There are 2 of the 8 that have exactly 2 heads; so the probability of exactly two coins landing on heads is 2/8 or 1/4.

Define the four different sampling techniques?

1) Simple random sampling

2) Systematic sampling

3) Stratified sampling

4) Cluster sampling

5) Probability proportional to size sampling

6) Matched random sampling

7) Quota sampling

8) Convenience sampling

9) Line-intercept sampling

10) Panel sampling

What are the odds of getting heads twice in a row with a coin?

"The same as getting heads the first time, a one-in-two chance."

No It's Not! "One in two chance" means 50/50 chance. The odds of getting 2 in a row are not 50/50. I don't know what it is because that's how I arrived at this question. But if my friend were to bet me that he could get to heads in a row.... I'd take him up on that bet everytime, saying that he couldn't. He'd win a couple times here and there but in the end would lose money. It's not 50/50.

Find two consecutive integers whose sum is 45?

First suppose the two consecutive integers bex and x+1then according to the situationx+x+1=452x+1=452x=45-12x=44x=44/2x=22now our assumed consecutive integers were x and x+1 , Now put x in itx=45x+1=45+1 =46Thus the Solution set is {(45,46)}

What is the probability of drawing 3 hearts in a row?

The probability of drawing 3 hearts in a row in a 52 card deck is 13/52 * 12/51 * 11/50 = 1716/132600 = 0.01294, or approximately 1 chance in 77.

What is the probability of drawing a king and a diamond from a deck of playing cards?

The probability of drawing a king is 4:52
The probability of drawing a diamond is 13:52 (or 1:4)
The probability of drawing a king (0.07692...) then replacing that king into the deck then drawing a diamond is 0.019230769.

If you leave the king out, the probability will be slightly greater (4/52) * (13/51)

Unless the king you left out of the deck was a king of diamonds, in which case, the probability would be (4/52) * (12/51)

Why event probabilities range from zero to one?

The probability of an event is the number of outcomes that are favourable to the outcome divided by the total number of outcomes. For continuous variables, it is the proportion of the outcome space, but the same argument applies.

The number of favourable outcomes must be non negative and so the probability has to be greater than or equal to 0. Also, the number of favourable outcomes can, at most, be as large as the total number so the probability must be less than or equal to 1.

If you flipped a coin ten times what is the probability you would get five heads?

We need to calculate two things:
How many possible possible series of 10 coin flips are there? As we flip 10 times and each time we can have either heads or tails we have 2 by the power of ten possibilities, or a total of 1024 unique possible series.

Now, how many of those series have exactly five heads and five tails? Lets assume we have ten "pre filipped" coins at hand - 5 tails and 5 heads. How many possible combinations are there. Well, if they were all different, you would have 10! (10 factorial = 10*9*8*7*6*5*4*3*2*1) possibilities.
How ever, the 5 heads are identical and so are the 5 tails, so if I interchange the locations of two coins that are both heads for example I still get the exact same series. There are 5! possible heads combinations, and 5! tails combinations.
Thus, the total number of unique combinations is 10!/(5!*5!) which happens to be 252.

So, out of 1024 possible series, 252 contain exactly 5 heads.
The probability thus is 252/1024=0.24609375 (roughly 25%)

Two events A and B with probability 0.5 and 0.7 respectively have joint probability of 0.4 The probability that neither A nor B happens is?

Let me denote -A as the event that A does not happen. So we want Pr[-(A and B)]

Now, the event that neither A nor B occurs is the opposite of either A occurring, or B occurring or both occurring.

So Pr[-(A and B)] = 1 - Pr(A or B)= 1 - [Pr(A) + Pr(B) - Pr(A and B)] (since A+B is double counted)= 1 - (0.5 + 0.7 - 0.4)= 1 - 0.8= 0.2

If the overall percentage of success in exam is 60 .what is probability that ouit of group of 4 students.at least one has passed?

This is a Binomial Probability Distribution; n=4, p=0.6. The probability of at least 1 passed is equal to the probability of 1-none passed; so x=0. The probability of x=0 (with n=4, p=0.6) is 0.0256. So, the probability of at least 1 passed is 1-0.0256 or 0.9744.